# M3-Net: A Cost-Effective Graph-Free MLP-Based Model for Traffic Prediction

### v1

@@ Abstract.
Achieving accurate traffic prediction is a fundamental but crucial task in the development of current intelligent transportation systems. The main challenges of the existing deep learning approaches are that they either depend on a complete traffic network structure or require intricate model designs to capture complex spatio-temporal dependencies. These limitations pose significant challenges for the efficient deployment and operation of deep learning models on large-scale datasets. To address these challenges, we propose a cost-effective graph-free Multilayer Perceptron (MLP) based model M3-Net for traffic prediction. Our proposed model not only employs time series and spatio-temporal embeddings for efficient feature processing but also first introduces a novel MLP-Mixer architecture with a mixture of experts (MoE) mechanism. Extensive experiments conducted on multiple real datasets demonstrate the superiority of the proposed model in terms of prediction performance and lightweight deployment.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 3.3. Ablation Study
From Table 3, we can find that our complete model M3-Net outperforms all the ablation variants. The comparison of w/o MOE can demonstrate that capturing multi-scale spatio-temporal dependencies of traffic flow in road networks can be beneficial for predictive performance. The variant of w/o Spatial MLP can illustrate that the spatial MLP can capture the spatial correlations in historical traffic sequences for more accurate prediction. To further investigate the effectiveness of adaptive group matrix, we compare M3-Net with w/o Grouping Matrix. This comparison indicates that capturing spatial homogeneity is crucial for prediction performance. If the adaptive grouping matrix learning module is removed, it may not only prevent the learning of spatial homogeneity features but also lead to performance degradation due to overfitting.
@@ 4. CONCLUSION
In this paper, we propose M3-Net, a lightweight and effective graph-free Multilayer Perceptron model to addresses the critical need for accurate traffic prediction in intelligent transportation systems. The innovative MLP-Mixer architecture, enhanced with a mixture of experts mechanism, allows for the joint learning of spatial and channel dimensions in traffic flow data. The incorporation of adaptive grouped matrices within the spatial MLP module effectively reduces the complexity of spatial learning, facilitating a more streamlined and efficient model. The extensive experiments conducted on various real-world datasets validate the effectiveness of M3-Net, demonstrating its superior performance in traffic prediction while maintaining a lightweight and deployable structure.

### v2

@@ Abstract.
Achieving accurate traffic prediction is a fundamental but crucial task in the development of current intelligent transportation systems. The main challenges of the existing deep learning approaches are that they either depend on a complete traffic network structure or require intricate model designs to capture complex spatio-temporal dependencies. These limitations pose significant challenges for the efficient deployment and operation of deep learning models on large-scale datasets. To address these challenges, we propose a cost-effective graph-free Multilayer Perceptron (MLP) based model M3-Net for traffic prediction. Extensive experiments conducted on multiple real datasets demonstrate the superiority of the proposed model in terms of prediction performance and lightweight deployment.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Korea.booktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreaisbn: 979-8-4007-2040-6/2025/11doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Spatial-temporal systemsisbn: 978-1-4503-XXXX-X/2018/06
@@ 3.3. Ablation Study
From Table 3, we can find that our complete model M3-Net outperforms all the ablation variants. The comparison of w/o MOE can demonstrate that capturing multi-scale spatio-temporal dependencies of traffic flow in road networks can be beneficial for predictive performance. The variant of w/o Spatial MLP can illustrate that the spatial MLP can capture the spatial correlations in historical traffic sequences for more accurate prediction. To further investigate the effectiveness of adaptive group matrix, we compare M3-Net with w/o Grouping Matrix. This comparison indicates that capturing spatial homogeneity is crucial for prediction performance.
@@ 4. CONCLUSION
In this paper, we propose M3-Net, a lightweight and effective graph-free Multilayer Perceptron model to addresses the critical need for accurate traffic prediction in intelligent transportation systems. The innovative MLP-Mixer architecture, enhanced with a mixture of experts mechanism, allows for the joint learning of spatial and channel dimensions in traffic flow data. The incorporation of adaptive grouped matrices within the spatial MLP module effectively reduces the complexity of spatial learning, facilitating a more streamlined and efficient model. The extensive experiments conducted on various real-world datasets validate the effectiveness of M3-Net, demonstrating its superior performance in traffic prediction while maintaining a lightweight and deployable structure. Future work will explore further enhancements to the M3-Net architecture and its applicability to other domains within transportation and urban planning.